OpenStudy (anonymous):

A rectangle has length 4x–1 and a width 2x+1. 1. a) Show how to find the perimeter of the rectangle. b) Simplify the expression. 2. a) Write an algebraic expression for the area of the rectangle. b) Simplify the expression. 3.If the area of another rectangle is x^2+5x+4, find algebraic expressions representing possible lengths and widths of the rectangle.

OpenStudy (karatechopper):

ok! simple!! so u have the length and width in front of u! soo this is like a reg perimeter problem!

OpenStudy (anonymous):

1) Multiply both sides by 2 and add them together. 4x - 2 + 4x + 2 = 8x 2) (4x - 1)(2x + 1) 8x^2 + 2x - 1 3) x^2 + 5x + 4 (x + 4)(x + 1)

OpenStudy (karatechopper):

2L+2W=P

OpenStudy (anonymous):

Basically you would add use the rectangle formula and do the following: perimeter=2l+2w perimeter=2(4x-1)+2(2x+1) perimeter=? (that's for you to answer =) ) The area would be the same except with a different formula. The area of a rectangle is length times width sooo: (4x-1)(2x+1)=A And then distribute from there. to solve for the last one, you would need to factor. Sooo: x^2+5x+4=0 As far as I know, that's unfactorable, so you're going to have to substitute in values manually. Have fun =)

OpenStudy (anonymous):

thanks everyone, you all helped alot ! xoxo :)